OpenStudy (anonymous):

Graph the following equations and inequalities on a coordinate plane, shading the solution set where necessary. y= 4x-2

OpenStudy (anonymous):

i think this is what the graph should look like and there should be no shading since there is an = sign. is that correct?

jimthompson5910 (jim_thompson5910):

y= 4x-2 is an equation due to the equal sign When graphing equations, you graph a curve or a line. No shading is done. So you are correct with that aspect. However, the graph is incorrect because notice how y = 4x-2 has a positive slope (the slope is +4), but the graph you posted has a negative slope. It goes downhill as you go from left to right.

jimthompson5910 (jim_thompson5910):

So there is no way that the equation and the graph match

OpenStudy (triciaal):

"Graph the following equations and inequalities" I think you should have an inequality in the original post. "shading the solution set where necessary." additional support that you forgot part of the original post. please verify the question

OpenStudy (anonymous):

@jim_thompson5910 can you show me a picture of what the graph should look like

OpenStudy (anonymous):

@triciaal i think what your saying is the way i have worded my question is confusing. is that correct? i should have put graph the equation y= 4x-2 and then also add shade solution when necessary.

jimthompson5910 (jim_thompson5910):

when x = 0, we plug this into the equation to get.. y = 4x - 2 y = 4*0 - 2 ... replace every x with 0 y = 0 - 2 y = -2 So when x = 0, y = -2. We have the ordered pair (0,-2) What happens when x = 1? What is the value of y?

OpenStudy (anonymous):

y=4x-2 y=4*1-2 y=2 so (1,2) ?

jimthompson5910 (jim_thompson5910):

correct, you now know that the points (0,-2) and (1,2) are on this line

jimthompson5910 (jim_thompson5910):

the min number of points you need to make a straight line is 2 points so plot them and draw a straight line through them to graph y = 4x-2

jimthompson5910 (jim_thompson5910):

you will notice that it slopes upward as you read from left to right

OpenStudy (triciaal):

" i think what your saying is the way i have worded my question is confusing. is that correct? i should have put graph the equation y= 4x-2 and then also add shade solution when necessary." this is the same as the posted question. I am saying you need to have 2 situations to find the intersection. All you have is one line. What are you comparing it with? When you compare it with the other situation then you can shade the intersection. that will be the solution.

OpenStudy (triciaal):

now that @jim_thompson5910 has shown you how to graph the line y = 4x - 2 you need another equation or an inequality to compare it with